# Mind the Gap: a Spectral Analysis of Rank Collapse and Signal Propagation in Transformers

### v1

@@ Abstract
Attention layers are the core component of transformers, the current state-of-the-art neural network architecture. However, softmax-based attention causes transformers to be more challenging to train. Even at initialisation, the propagation of signals and gradients through the random network can be pathological, resulting in known issues such as (i) vanishing/exploding gradients and (ii) rank collapse, i.e. when all tokens converge to a single representation with depth. This paper examines signal propagation in attention-only transformers from a random matrix perspective, illuminating the origin of such issues, as well as unveiling a new phenomenon—(iii) rank collapse in width. Modelling softmax-based attention at initialisation with Random Markov matrices, our theoretical analysis reveals that a spectral gap between the two largest singular values of the attention matrix causes (iii), which, in turn, exacerbates (i) and (ii). Building on this insight, we propose a novel, yet simple, practical solution to resolve rank collapse in width by removing the spectral gap. Moreover, we validate our findings and discuss the training benefits of the proposed fix through experiments [^2] Our code is available at https://shorturl.at/0zk8q. that also motivate a revision of some of the default parameter scaling. Our attention model accurately describes the standard key-query attention in a single-layer transformer, making this work a significant first step towards a better understanding of the initialisation dynamics in the multi-layer case.
@@ 1 Introduction
The purpose of this work is to analyse signal propagation in softmax-based attention layers at initialisation, i.e. with randomly initialised model parameters. While the issues of rank collapse (in depth) and vanishing/exploding gradients have been previously identified in transformers at initialisation [DCL21, NAB+22], our work formalises these findings and uncovers an additional phenomenon—rank collapse in width—due to the use of softmax in the self-attention mechanism. Rank collapse in width has not been identified in the existing literature nor been recognised as a catalyst for rank collapse along depth. By leveraging spectral properties of the random matrices formed by the model’s parameters, we reveal the emergence of a spectral gap between the two largest singular values of the attention matrix, which drives rank collapse in width and further accelerates rank collapse in depth. Moreover, we propose a provably effective remedy for the spectral gap, a solution that naturally arises when the problem is viewed through a spectral lens. To the best of our knowledge, a spectral analysis of signal propagation has yet not been undertaken in the context of transformers.
@@
Let us consider the eigenvalues of an attention matrix. Since the rows sum to 1, there is an eigenvalue of 1 corresponding to the all-ones vector. Under certain conditions, the other eigenvalues shrink in size as the matrix dimension increases, resulting in a widening gap between the largest eigenvalue (which is 1) and the diminishing bulk of eigenvalues; see Figure 1. The successive multiplication of such matrices at each layer increasingly favours a specific direction—the one aligned with the dominant eigenvector of the attention matrix—over the others. This leads to a distortion in the geometry of the input training data, exemplified by the phenomenon of rank collapse. A natural solution is then to project out this troublesome direction from all attention matrices to enable a more balanced signal propagation. This intuitive idea is central to our rigorous analysis of a simple transformer, from which we draw insights to introduce a slightly modified attention layer that proves advantageous even when incorporated into more complex architectures.
Spectra of random matrices.
@@
converge, in many interesting cases, to deterministic distributions $\mu$ and $\nu$, known as the limiting eigenvalue/singular value distribution of $\mathbf{M}_{n}$. Additionally, the largest eigenvalue/singular value of random matrices is often studied in its own right. Our analysis builds on several established results concerning both the limiting distribution of the eigenvalues/singular values (the “bulk” of the spectrum) and the behaviour of the largest eigenvalue/singular value (the “edge” of the spectrum). In particular, we focus on two classes of random matrices: Gaussian and Markov, that respectively model the value and attention matrices in our transformer model ((3)) at initialisation.
Free probability.
@@ 1.1 Attention at initialisation
We model the attention mechanism at initialisation by a random matrix $\mathbf{A}_{\ell}$ with non-negative entries $(\mathbf{A}_{\ell})_{i,j}\geq 0$ and normalised rows, i.e. $\sum_{j}(\mathbf{A}_{\ell})_{i,j}=1$, as if it were generated by a row-wise application of softmax. As we will demonstrate, this model functions as a helpful abstraction that offers insights into the causes of rank collapse. More specifically, we consider $\mathbf{A}_{\ell}$ to be a Random Markov matrix, as defined in [BCC11].
Definition 1.1 (Random Markov matrix).
@@
We call $\mathbf{A}$ a Random Markov matrix.
(a) Random Markov matrix of Definition 1.1
(b) Standard key-query attention at layer one
Figure 1: As the size $T$ of a Random Markov matrix grows, its eigenvalues form a circular bulk of radius $O(T^{-1/2})$ in the complex plane, except for the largest eigenvalue which remains 1 (the black dot). Proposition 1 demonstrates that applying the conventional key-query attention mechanism to orthonormal input tokens yields a Random Markov matrix. In practice, $T$ does not need to be too large for the limiting behaviour to appear, as shown above.
Note that not all random Markov matrix ensembles satisfy the conditions of Definition 1.1 which we colloquially refer to as Random Markov matrix. Remarkably, the standard key-query dot product attention matrix in the first layer of a transformer is a Random Markov matrix, as specified in the following.
Proposition 1.
Let $\mathbf{X}_{0}\in\mathbb{R}^{T\times d}$ have orthonormal rows, i.e. $\mathbf{X}_{0}\mathbf{X}^{\top}_{0}=\mathbf{I}$. Let
$\mathbf{A}_{1}(\mathbf{X}_{0})\coloneqq\text{softmax}\bigg{(}\frac{\mathbf{X}_{0}\mathbf{W}_{1}^{Q}{\mathbf{W}^{K}_{1}}^{\top}\mathbf{X}^{\top}_{0}}{\sqrt{d_{qk}}}\bigg{)},$
where $\mathbf{W}^{Q}_{1},\mathbf{W}^{K}_{1}\in\mathbb{R}^{d\times d_{qk}}$ have i.i.d. $\mathcal{N}(0,\sigma^{2}_{qk})$ entries. Then $\mathbf{A}_{1}(\mathbf{X}_{0})$ is a Random Markov matrix as in Definition 1.1 when $d_{qk}$ is large  [^3] Standard transformer implementations set $d_{qk}=d$, as detailed in the original paper by [VSP+17]. Therefore, assuming large $d_{qk}$ is not restrictive since we analyse the network in the large $d$ regime.. In particular, under Xavier or He scaling of $\mathbf{W}_{1}^{Q},\mathbf{W}_{1}^{K}$ ([GB10, HZRS15], or any scaling such that $\sigma_{qk}^{2}\to 0$ as $d_{qk}$ increases), the resulting $\mathbf{A}_{1}(\mathbf{X}_{0})$ is degenerate, i.e. the “uniform attention” $\frac{1}{T}\mathbf{1}_{T\times T}$.
The latter part of the above proposition was shown in [NAB+22, Lemma A.7] to justify their assumption of uniform attention, under which they demonstrate rank collapse in depth. In contrast, our analysis of the rank collapse is also valid for non-degenerate attention matrices. Based on the above proposition, much of our analysis is relevant to single-layer transformers using the standard key-query attention mechanism. However, as we will demonstrate in section 3, the case of multi-layer real-world transformers presents additional complexities, warranting further investigation.
It is shown in [BCC11] that the “bulk” of the limiting eigenvalue/singular value distribution of a Random Markov matrix matches (up to a scaling) that of an i.i.d. Gaussian matrix. Namely, if $\mathbf{A}$ is as in Definition 1.1 with variance $\sigma^{2}$, then $\sqrt{T}\mathbf{A}$ has the same bulk density as a Gaussian matrix with i.i.d. $\mathcal{N}(0,\frac{\sigma^{2}}{T})$ entries. Nonetheless, the “edge” of the spectrum of $\sqrt{T}\mathbf{A}$ behaves quite differently from the i.i.d. case. While the largest eigenvalue/singular value of an i.i.d. matrix is almost surely right at the boundary of its bulk, for Random Markov matrices there is a gap between the edge and the bulk given by Theorem 2. Without loss of generality, we formulate the theorem for Random Markov matrices with mean $1$.
Theorem 2 (Spectral gap in Random Markov matrices, [BCC11]).
Let $\mathbf{A}\in\mathbb{R}^{T\times T}$ be a Random Markov matrix whose underlying distribution has variance $\sigma_{A}^{2}$. Then, $\lambda_{1}(\mathbf{A})=1$ and almost surely,
[…]
1.2 Related work
Rank collapse in transformers was first explored in [DCL21], where the authors show that the output of an attention-only transformer converges exponentially with depth to a single representation across tokens. A connection between rank collapse and vanishing gradients is made in [NAB+22]. By assuming uniform attention, the work of [NAB+22] essentially reduces to proving rank collapse in depth, based on the assumption that rank collapse in width has already occurred (as opposed to showing why this premise holds). Our spectral analysis builds upon previous work that analysed the spectra of large random neural networks to better understand and stabilise initial training dynamics, such as [PSG18] for fully-connected networks and [XBSD+18] for convolutional networks.
It is also worth mentioning a line of research on possible alternatives to the softmax self-attention mechanism. [HBSDN20] speculate the advantage of ReLU and identity over softmax in training accuracy. [HMZ+22] proposes an ad hoc initialisation scheme tailored to prevent the token-wise covariance kernel from losing rank through layers. Besides, several practical works attempt to replace softmax-based key-query attention with faster options that surpass the so-called quadratic bottleneck, e.g. [PPY+20, CLD+20, KVPF20].
1.3 Organisation of the paper
In section 2, we first introduce our model and reexamine the phenomena of rank collapse and exploding gradients, demonstrating that both occur with increasing depth in our transformer model at initialisation (Props. 3 and 4). Importantly, we diagnose for the first time an additional form of rank-collapse with increasing width, that we call rank collapse in width. We formulate its exact rate of decay in the context length as well as identify its root cause as being the spectral gap in softmax-based attention. Next, we introduce a modified attention mechanism that is specifically designed to have no spectral gap. We prove that this modification simultaneously resolves rank collapse in width, thus mitigating rank collapse in depth, and exploding gradients (Props. 5 and 6), thereby confirming the role of the spectral gap in such issues. Lastly, we study the spectra of the covariance kernel and the input-output Jacobian in our modified model (Props. 7 and 8) and discuss the possibility of further stabilising early training dynamics by tuning the initial distributions.
In section 3, we validate our findings, providing empirical evidence of rank collapse in both width and depth, as well as exploding gradients. We put to test our “remove the gap” solution across a range of architectures featuring LayerNorm and skip connections, discussing its possible training benefits. Finally, we present experiments that challenge the soundness of certain default scaling choices, such as Xavier initialization for the keys and queries, suggesting that they may require further revision in practice.
2 Theoretical Results
@@
We study as our model a deep attention-only single-head transformer at initialisation, where at each layer $\ell$, the signal is transformed as $\mathbf{X}_{\ell}=\mathbf{A}_{\ell}\mathbf{X}_{\ell-1}\mathbf{W}^{V}_{\ell}$. The input signal $\mathbf{X}_{0}\in\mathbb{R}^{T\times d}$ has $T$ tokens of embedding dimension $d$, with a fixed ratio $\gamma\coloneqq\frac{T}{d}\leq 1$. For a network of depth $L$, the input-output relationship is thus given by
@@
The value matrices $\mathbf{W}^{V}_{\ell}\in\mathbb{R}^{d\times d}$ are initialised independently with i.i.d. $\mathcal{N}(0,1)$ entries and the attention matrices $\mathbf{A}_{\ell}\in\mathbb{R}^{T\times T}$ are independent Random Markov matrices with $\sigma^{2}_{A}=1$.

### v2

# Mind the Gap: a Spectral Analysis of Rank Collapse and Signal Propagation in Attention Layers

@@ Abstract
Attention layers are the core component of transformers, the current state-of-the-art neural network architecture. Alternatives to softmax-based attention are being explored due to its tendency to hinder effective information flow. Even at initialisation, it remains poorly understood why the propagation of signals and gradients through these random networks can be pathological, resulting in issues known as (i) vanishing/exploding gradients and (ii) rank collapse in depth, i.e. when all tokens converge to a single representation along layers. While rank collapse in depth naturally arises from repeated matrix multiplications—a common pattern across various architectures—we identify an additional and previously unknown challenge unique to softmax attention layers: (iii) rank collapse in width, which occurs as the context length increases. Using Random Matrix Theory, we conduct a rigorous analysis that uncovers a spectral gap between the two largest singular values of the attention matrix as the cause of (iii), which in turn exacerbates (i) and (ii). Building on this insight, we propose a novel yet simple practical solution to mitigate rank collapse in width by removing the outlier eigenvalue(s). Our theoretical framework offers a fresh perspective on recent practical studies, such as [YDX+24, AGW23], whose ad hoc solutions can now be interpreted as implicit efforts to address the spectral gap issue. This work provides valuable theoretical support for ongoing large-scale empirical research, bringing theory and practice one step closer in the understanding of transformers. [^2] Our code is available at https://shorturl.at/0zk8q.
@@ 1 Introduction
The purpose of this work is to analyse signal propagation in softmax-based attention layers at initialisation, i.e. with randomly initialised model parameters. While the issues of rank collapse (in depth) and vanishing/exploding gradients have been previously identified in transformers at initialisation [DCL21, NAB+22], our work extends these findings and uncovers an additional phenomenon—rank collapse in width—due to the use of softmax in the self-attention mechanism. By width, we specifically refer to the context length, which, despite its increasing scale in modern attention layers, seems to have been overlooked in favour of depth in previous analyses of rank collapse in transformers. Indeed, rank collapse in width, which is unique to attention layers, has not been identified in the existing literature nor recognised as a catalyst for rank collapse along the depth. In contrast, rank collapse in depth is not specific to transformer-like networks, as it typically results from successive matrix multiplications, a shared feature of various deep neural network architectures. Investigating the spectra of the random matrices formed by the model’s parameters, we detect a spectral gap between the two largest eigenvalues of the attention matrix, which drives rank collapse in width and further accelerates rank collapse in depth. Moreover, we propose a provably effective remedy for the spectral gap, a solution that naturally arises when the problem is viewed through a spectral lens. To the best of our knowledge, this is the first study to examine the attention mechanism in the large context length regime, as well as the first to identify rank collapse in width and its effect on other oversmoothing phenomena.
@@
Let us consider the eigenvalues of a softmax-based attention matrix. Since the rows sum to unity, there is an eigenvalue of $1$ corresponding to the eigenvector of all-ones. Under certain conditions, the other eigenvalues shrink in size as the matrix dimension increases, resulting in a widening gap between the largest eigenvalue (i.e. $1$) and the diminishing bulk of eigenvalues; see Figure 1. The successive multiplication of such matrices at each layer increasingly favours a specific direction—the one aligned with the dominant eigenvector of the attention matrix—over the others. This leads to a distortion in the geometry of the input training data, exemplified by the phenomenon of rank collapse. An intuitive solution is to project out this troublesome direction from all attention matrices, allowing for more balanced signal propagation. These ideas lie at the core of our rigorous analysis of attention layers, from which we derive a slightly modified attention mechanism that proves advantageous in practice. Remarkably, the spectral gap observed in softmax attention matrices also appears in alternative attention mechanisms suggested in the literature, making our analysis relevant to these variants as well; see Figure 1.
1.1 Related work
Our analysis builds upon previous work that studied the spectra of large random neural networks to better understand and stabilise initial training dynamics, such as [PSG18] for fully-connected networks and [XBSD+18] for convolutional networks.
The rank collapse in transformers was first explored in [DCL21], where the authors show that the output of an attention-only transformer converges exponentially with depth to a single representation across tokens. [NAB+22] make a connection between rank collapse and vanishing gradients by assuming “uniform attention” $\mathbf{A}=\frac{1}{T}\mathbf{1}_{T\times T}$, essentially proving rank collapse in depth based on the assumption that rank collapse in width has already occurred (as opposed to showing why this premise holds).
There is also a growing literature on the shortcomings of softmax self-attention mechanism and possible alternatives. Some suggest the usage of other well-known activations such as identity, ReLU, and sigmoid [HBSDN20, WLGK23, RDD+24], while others propose modifications to the attention block [HMZ+22, WZCW22, SGX+22, AGW23, DBK24, YDX+24]. Based on our analysis, we can offer a fresh perspective on these attention variants and reflect on their effect on rank collapse. Indeed, they all fall (at the exception of the identity) within the class of attention matrices presenting a spectral gap, therefore suffering from rank collapse, with slightly adjusted rates according to the location of the outlier in the spectrum; see Figure 1.
Interestingly, our work and the fix we propose find echos in recent promising practical studies. For instance, the Differential Transformer [YDX+24] computes the attention matrix as a difference between two softmax matrices and has shown significant improvements from this adjustment, e.g. finer quantisation, better key information retrieval or hallucination mitigation. In light of our findings, this adjustment can be understood as removing the troublesome outlier from the spectrum of the attention matrix that naturally arises when using softmax. Even closer, the work of [AGW23] directly implements our fix and shows improved behaviour on image segmentation tasks. When our work is of theoretical nature and concerned with the initialisation dynamics, it is interesting to see how these concepts are playing out in practical applications.
1.2 Organisation of the paper
In Section 2, we first review the fundamentals of random matrix spectra before exploring the properties of standard softmax key-query attention matrices with isotropic input. We prove the existence of a spectral gap for such attention matrices and use this result to demonstrate rank collapse in width. Next, we introduce a modified attention mechanism designed to eliminate the spectral gap and show that this modification effectively resolves rank collapse in width.
In Section 3, we extend our analysis from a single attention layer to a deep attention-only transformer. Using a certain class of random Markov matrices to model attention, we precisely determine the rate at which rank collapse occurs as a function of width and depth (Proposition 4), thereby revealing for the first time how the two interact. Furthermore, we prove that the removal of the spectral gap cures rank collapse in width and mitigates the exploding gradients (Propositions 6 and 7).
In Section 4, we validate our findings, providing empirical evidence of rank collapse in both width and depth in commercial models such as BERT family. We put to test our “remove the gap” solution across a range of architectures featuring LayerNorm and skip connections and show its effectiveness in resolving stable rank collapse and slowing down gradient explosion. Finally, we observe the emergence of additional outliers in deep transformers and conduct experiments to study the potential benefits of removing all outliers rather than the dominant one.
(a) Softmax attention [VSP+17]
(b) i.i.d. Markov
(c) ReLU attention [WLGK23]
(d) Sigmoid attention [RDD+24]
Figure 1: Theorem 1 demonstrates that applying the conventional softmax key-query attention mechanism to orthonormal input tokens yields a spectral gap in its spectrum (a). Due to their non-zero mean, ReLU (c) and sigmoid attention (d) mechanisms also exhibit a spectral gap—although of a different size. As the size $T$ of an i.i.d. Markov matrix (Definition 3.1) grows, its eigenvalues form a circular bulk of radius $O(T^{-1/2})$ in the complex plane, except for the largest eigenvalue which remains 1 (the black dot in (b)). In practice, $T$ does not need to be too large for the limiting behaviour to appear, as shown above.
2 Attention layers and rank collapse in width
2.1 Spectra of random matrices
@@
converge, in many interesting cases, to deterministic distributions $\mu$ and $\nu$, known as the limiting eigenvalue/singular value distribution of $\mathbf{M}_{n}$. Additionally, the largest eigenvalue/singular value of random matrices is often studied in its own right. Our analysis builds on cutting-edge results concerning both the limiting distribution of the eigenvalues/singular values (the “bulk” of the spectrum) and the behaviour of the largest eigenvalue/singular value (the “edge” of the spectrum). In particular, we focus on two classes of random matrices: Gaussian and Markov, that respectively model the value and attention matrices in our transformer model (11) at initialisation.
@@ 2.2 Attention at initialisation
Let $\mathbf{X}_{0}\in\mathbb{R}^{T\times d}$ be an input sequence of context length $T$, i.e. formed of $T$ tokens. Each token is embedded as a $d$-dimensional vector, with a fixed ratio $\gamma\coloneqq\frac{T}{d}\leq 1$. The standard softmax key-query attention matrix is obtained by
$\mathbf{A}(\mathbf{X}_{0})\coloneqq\mathrm{softmax}\bigg{(}\frac{\mathbf{X}_{0}\mathbf{W}^{Q}{\mathbf{W}^{K}}^{\top}\mathbf{X}^{\top}_{0}}{\sqrt{d_{qk}}}\bigg{)},$ (1)
where $\mathbf{W}^{Q},\mathbf{W}^{K}\in\mathbb{R}^{d\times d_{qk}}$ have i.i.d. $\mathcal{N}(0,\sigma^{2}_{qk})$ entries.
The above attention matrix exhibits a spectral gap, meaning its largest eigenvalue (or singular value) differs significantly in magnitude from the second largest. Moreover, this gap widens as the context length increases. In fact, we precisely quantify the rates at which both eigenvalues evolve, as follows.
Theorem 1 (Spectral gap in softmax attention).
Let the input sequence $\mathbf{X}_{0}$ have orthonormal rows, i.e. $\mathbf{X}_{0}\mathbf{X}^{\top}_{0}=\mathbf{I}$. Then, $\lambda_{1}(\mathbf{A}(\mathbf{X}_{0}))=1$ and almost surely,
$\lim_{T\to\infty}s_{1}(\mathbf{A}(\mathbf{X}_{0}))=1,$ (2)
and
$\lim_{T\to\infty}s_{2}(\sqrt{T}\mathbf{A}(\mathbf{X}_{0}))=2\sigma,$ (3)
while
[…]
where $\sigma$ depends only on $\sigma_{qk}$.
[…]
Figure 2: The standard key-query attention matrices of Equation (1) form a subclass of i.i.d. Markov matrices of Definition 3.1, of which the uniform attention of [NAB+22] is one trivial special case. We use i.i.d. Markov matrices to build our general multi-layer theory in section 3.
With no further assumptions, we exactly analyse one attention layer at initialisation, where the output is given by
$\mathbf{X}=\mathbf{A}(\mathbf{X}_{0})\mathbf{W}^{V},$
and the value matrix $\mathbf{W}^{V}\in\mathbb{R}^{d\times d}$ is initialised independently with i.i.d. $\mathcal{N}(0,1)$ entries  [^3] See Remark Remark for a discussion on scaling the value matrix..
2.3 Rank collapse in width
Theorem 1 reveals that the attention matrix becomes effectively rank-one for large context length $T$. Naturally, any definition of “rank collapse” relies on a proxy for discrete rank, as the random matrices in question are almost surely full-rank, making it uninformative to refer to their actual rank. For example, [DCL21] consider the “one-infinity norm” of the residual (the difference between a matrix and its best approximation of the form $\mathbf{1}\mathbf{x}^{\top}$), defined as $\sqrt{\lVert\mathrm{res}(\mathbf{M})\rVert_{1}\lVert\mathrm{res}(\mathbf{M})\rVert_{\infty}}$, while [NAB+22] use $\sum_{i,j}(\mathbf{M}\mathbf{M}^{\top})_{i,j}$, which is maximised when all rows of $\mathbf{M}$ are identical. We choose stable rank as our preferred proxy due to its clear geometrical interpretation and simple definition in terms of singular values. It is defined for any non-zero $\mathbf{M}\in\mathbb{R}^{m\times n}$ as
The following theorem establishes that the stable rank of the covariance kernel
$\boldsymbol{\Sigma}\coloneqq\mathbf{X}\mathbf{X}^{\top}$
collapses to $1$ as the width $T$ increases.
Theorem 2 (Rank collapse in width).
Assume $\boldsymbol{\Sigma}_{0}=\mathbf{I}$. Then, with overwhelming probability [^4] A sequence of events $E_{n}$ holds with overwhelming probability if, for every $A>0$, $\mathbb{P}(E_{n})\geq 1-C_{A}n^{-A}$, for some constant $C_{A}$, as defined in [Tao12].,
$\lim_{T\to\infty}\mathrm{sr}(\mathbf{\Sigma})=1.$ (6)
Moreover, the convergence happens at a polynomial rate, i.e. $\big{|}\mathrm{sr}(\boldsymbol{\Sigma})-1\big{|}=O(T^{-3})$.
This theorem constitutes the primary contribution of this paper. It uncovers a previously unknown phenomenon that we refer to as rank collapse in width—or the convergence of tokens’ representations into a single point as the context length increases within a single attention layer. It should not be mistaken for rank collapse in depth, which is partly a result of repeated random matrix multiplications. In contrast, rank collapse in width is unique to attention layers, making this discovery particularly significant for the theory of transformers. In the next section, we will explore how removing the spectral gap of Theorem 1 helps alleviate rank collapse in width and restore a more balanced flow of information.
2.4 Attention without the gap
Assume $\mathbf{X}_{0}\mathbf{X}_{0}^{\top}=\mathbf{I}$ and let us write $\mathbf{A}\coloneqq\mathbf{A}(\mathbf{X}_{0})$ to lighten the notation. The attention matrix $\mathbf{A}\in\mathbb{R}^{T\times T}$ can be written as
where $\mathbf{1}_{T\times T}\coloneqq(1,\cdots,1)^{\top}(1,\cdots,1)$ is the all-ones matrix. Therefore, $\mathbf{A}$ is a rank-one perturbation of $\mathbf{A}^{\perp}$, whose spectral radius is $\lambda_{1}(\mathbf{A}^{\perp})=O(T^{-1/2})$ (see Lemma 2). Although the rank-one perturbation $\frac{1}{T}\mathbf{1}_{T\times T}$ cannot disturb the bulk of the spectrum, it causes the largest eigenvalue to “escape” from the bulk to $1$, creating a spectral gap.
In light of this, we can slightly modify the attention mechanism to eliminate the outlier—and thus the gap—simply by replacing $\mathbf{A}$ with $\mathbf{A}^{\perp}$ to get the following modified attention layer, i.e.
$\mathbf{X}^{\perp}=\mathbf{A}^{\perp}(\mathbf{X}_{0})\mathbf{W}^{V},$ (8)
where $\mathbf{A}^{\perp}(\mathbf{X}_{0})\coloneqq\mathbf{A}(\mathbf{X}_{0})-\frac{1}{T}\mathbf{1}_{T\times T}$. Note that this modification is applied only to the attention matrix (and not to the signal representation) and $\mathbf{X}^{\perp}$ serves as shorthand for the output of an attention layer whose $\mathbf{A}$’s is replaced with $\mathbf{A}^{\perp}$’s as in (8).
Since the modified attention exhibits no spectral gap (see Lemma 3), the stable rank of the covariance matrix $\boldsymbol{\Sigma}^{\perp}\coloneqq\mathbf{X}^{\perp}{\mathbf{X}^{\perp}}^{\top}$ no longer collapses to 1 but instead scales linearly in width, as detailed in Proposition 3 (cf. Theorem 2).
Proposition 3 (Resolved rank collapse in width).
Let $\mathbf{X}^{\perp}=\mathbf{A}^{\perp}(\mathbf{X}_{0})\mathbf{W}^{V}$ be the signal in our modified attention layer (8) and $\mathbf{\Sigma}^{\perp}\in\mathbb{R}^{T\times T}$ be its covariance matrix. Then, almost surely, the rank does not collapse, i.e., there exists a constant $C>0$ such that,
$\lim_{T\to\infty}\frac{\mathrm{sr}(\mathbf{\Sigma}^{\perp})}{T}=C.$ (9)
3 The interplay between width and depth
It is important to note that a spectral gap within the attention matrix can arise when the input to the softmax function consists of a matrix with vanishing elements. This can be achieved, for example, by initialising the key and query matrices following Xavier or He initialisation schemes, or under any scaling such that $\sigma_{qk}^{2}\to 0$ as $d_{qk}$ increases. In this scenario, the attention matrix asymptotically approaches the uniform attention matrix, i.e. the matrix of all entries equal to $1/T$. In this extreme case, the attention matrix becomes rank-one, the spectral gap is maximised and once the rank-one perturbation $\frac{1}{T}\mathbf{1}_{T\times T}$ is removed, the attention is reduced to zero—effectively preventing any signal propagation beyond the rank collapse.
It is, nonetheless, in this trivial case that the analysis of signal propagation along depth was conducted in [NAB+22]. In particular, their study demonstrates rank collapse in depth by implicitly assuming that rank collapse in width had already occurred. In contrast, our work presents a more refined analysis of rank collapse by considering the effects of both width and depth jointly, offering, to the best of our knowledge, the most advanced theoretical study to date on the initialisation of transformers.
@@ 3 The interplay between width and depth
We model the $\ell$-th layer attention mechanism at initialisation by a random matrix $\mathbf{A}_{\ell}$ with non-negative entries $(\mathbf{A}_{\ell})_{i,j}\geq 0$ and normalised rows, i.e. $\sum_{j}(\mathbf{A}_{\ell})_{i,j}=1$, as if it were generated by a row-wise application of softmax. As we will demonstrate, this model functions as a helpful abstraction that offers insights into the interaction of width and depth. More specifically, we consider $\mathbf{A}_{\ell}$ to be an i.i.d. Markov matrix, as defined in [BCC11].
Definition 3.1 (i.i.d. Markov matrix).
@@
We call $\mathbf{A}$ an i.i.d. Markov matrix.
Remark (Terminology).
This is a slight abuse of language since the entries of an “i.i.d. Markov matrix” are not necessarily independent random variables because of the normalisation constraint. Moreover, not all random Markov matrix ensembles satisfy the above definition. Remarkably, the standard key-query dot product attention matrix $\mathbf{A}(\mathbf{X}_{0})$ defined in (1) is an i.i.d. Markov matrix provided the input is isotropic $\mathbf{X}_{0}\mathbf{X}_{0}^{\top}=\mathbf{I}$. The uniform attention is another instance from this class of matrices; see Figure 2.
@@
We study as our model a deep stack of attention layers at initialisation, where at each layer $\ell$, the signal is transformed as $\mathbf{X}_{\ell}=\mathbf{A}_{\ell}\mathbf{X}_{\ell-1}\mathbf{W}^{V}_{\ell}$. The input signal $\mathbf{X}_{0}\in\mathbb{R}^{T\times d}$ has $T$ tokens of embedding dimension $d$, with a fixed ratio $\gamma\coloneqq\frac{T}{d}\leq 1$. For such a network of depth $L$, the input-output relationship is thus given by
@@
The value matrices $\mathbf{W}^{V}_{\ell}\in\mathbb{R}^{d\times d}$ are initialised independently with i.i.d. $\mathcal{N}(0,1)$ entries and the attention matrices $\mathbf{A}_{\ell}\in\mathbb{R}^{T\times T}$ are independent i.i.d. Markov matrices with $\sigma^{2}_{A}=1$.